# CheMFi: A Multifidelity Dataset of Quantum Chemical Properties of Diverse Molecules

### v1

@@ 2 Methods
To build the CheMFi dataset from the WS22 database, 15,000 geometries were sampled from the original 120,000 geometries for each of these molecules. These were evenly sampled across the 120,000 geometries to ensure even coverage of the chemical space. Further details of the sampling are mentioned in the supplementary material Section S1. An even sampling of the original dataset ensures that there are sufficient geometries from all conformations of the molecule. In order to verify that this form of sampling did in fact cover even chemical space, UMAPs 26 were studied and the results were satisfactory indicating the uniform coverage of the chemical space. These UMAPs and related discussion on the inferences are reported in the supplementary section S1. Once these geometries were sampled, they were used to perform point calculations for the QC properties. The complete workflow of the dataset generation process is pictorially depicted in Figure 1.
@@
E(SCF) (15000, 5)/hE ‘SCF’
[…]
@@
Table 1: List of properties available in the CheMFi dataset. The corresponding dimension(s) and units of the properties are also given with the npz file key. From the WS22 database 9, 10
For the 9 molecules taken from the WS22 database, after 15,000 geometries were evenly sampled, QC calculations were performed at the TD-DFT level of theory with the CAM-B3LYP functional. For each geometry, five fidelities were calculated. These fidelities are the basis set choice of increasing size. In increasing hierarchy of the fidelity, these are: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. In the rest of the document, for the most part, these are referred to by their short-hand, i.e., STO3G up to TZVP. The TightSCF keyword was employed to ensure energy convergence of the order of $10^{-9}$ a.u. for each calculation. Resolution of Identity approximation (RIJCOSX) was employed in order to speed up the excitation energy calculations. For any calculation, the maximum memory usage was limited to 2.0 GB. In practice, the ORCA calculations did not use this amount of memory. A total of 10 vertical excitation energies were calculated with each fidelity for each geometry.
All calculations were performed with the ORCA(5.0.1) quantum chemistry package 27. From these calculations, a diverse set of QC properties were extracted including information of the vertical excitation states such as energies and oscillator strengths. The list of available multifidelity properties is given in Table 1. The Cartesian coordinates and atomic numbers are taken from the WS22 database. The SCF energies are reported in Hartree units. The first 10 vertical excitation energies are provided in $\rm cm^{-1}$ with their corresponding oscillator strengths and transition dipole moments (in a.u.). The molecular dipole moments are also a property included in the CheMFi dataset with both the nuclear and electronic contributions being separately cataloged in atomic units (a.u.). The rotational spectrum data is also included in the form of Rotational Constants (in $\rm cm^{-1}$) and the total molecular dipole moments (in a.u.) aligned along rotational axes. Details of data hosting and storage are reported in Section 3.
3 Data Records
The various QC properties of the CheMFi dataset are stored in separate NumPy (v 1.26.4) npz files for each molecule. These npz files have a dictionary-like format allowing for each property to be accessed via its corresponding key denoted in Table 1. Each property itself is stored as a NumPy ndarray with the first dimension being 15,000 corresponding to the number of geometries. Thus, the QC properties can be accessed by querying the right ID. For example, the SCF energies can be accessed with the key ’SCF’ returning a NumPy ndarray of size $15,000\times 5$ where the second dimension of the array corresponds to the five fidelities used. An example script to accessing the QC properties is shown in Listing LABEL:lst_codeaccess.
The dataset itself is hosted on Zenodo at https://zenodo.org/records/11636903 with a detailed README file documenting the key aspects of the dataset 22. The README also provides information on how to access the different properties using Python. For the CheMFi dataset, the various scripts involved in generating the data, including ORCA input files and shell scripts to extract properties from the ORCA log files, are stored in the code repository that can be accessed at https://github.com/SM4DA/CheMFi. In addition to these scripts, the code repository also contains Python scripts to perform multifidelity benchmarks on this dataset. These can be launched using the CLI and are a handy tool in setting benchmarks for this dataset using current state of art multifidelity methods.
@@ 3 Data Records
9#note that Python arrays start with 0
@@
4 Technical Validation
To validate the CheMFi dataset, The MFML and o-MFML models prescribed in Ref. 15 were tested in predicting SCF energies and the first vertical excitation energies. The multifidelity models are built for different baseline fidelities, which refers to the cheapest fidelity included in the model. For example, a baseline fidelity $f_{b}=$631G implies that the multifidelity model is built up of the fidelities 631G, SVP, and TZVP 15.
In addition to benchmarking the multifidelity models on properties of individual molecules, the models are also tested on using all the molecules of the dataset. For this purpose, the SCF property of all molecules are used to train one single MFML and o-MFML model. This is then tested on predicting the SCF energies of all molecules.
While these two broad tests serve as a benchmark for multifidelity models on this dataset, the benchmarks of the other properties and molecules are not reported here. However, it is to be pointed out that the scripts provided can be readily used to generate benchmarks for these cases using standard ML methods such as learning curves. All learning curves are reported for a 10-run average, that is, for 10 random shuffling of the training set as directed in Ref. 15.
4.1 Single Molecule Benchmarks
(a) Preliminary analysis of multifidelity structure of SCF energies for the SMA molecule. The three different preliminary tests for the hierarchy are performed as prescribed in Ref. 13.
(b) Learning curves for MFML and o-MFML for the SCF property of SMA as recorded in the CheMFi database. The reference single-fidelity KRR is also shown by training on TZVP only. The Laplacian kernel was used with a kernel width of 200.0 and regularization of $10^{-10}$. The Global SLATM 28 molecular descriptors were used.
@@ 4.1 Single Molecule Benchmarks
The technical validation carried out individually for SMA and o-HBDI molecules is in line with the experimental set-up of Ref. 13. A total of 12,288 training samples were chosen to build the multifidelity models. 712 samples were set aside as a validation set for the o-MFML model from Ref. 15, and the remaining 2,000 samples were used as a test set. The accuracy of the models are gauged with mean absolute error (MAE) in the form of learning curves. Learning curves display the MAE with respect to increasing number of training samples, here, at the highest fidelity, that is, TZVP.
Figure 2 reports the multifidelity benchmarking process for SCF energies for SMA. The SLATM molecular descriptor 28 was used with a Laplacian kernel to perform kernel ridge regression (more details in Supplementary Information Section S2). First, the preliminary analyses as recommended by Ref. 13 are shown in Figure 2(a). These are recommended preliminary investigation of multifidelity data. The first of these is to study the distribution of the multifidelity data. The second analysis is the study of mean absolute differences between each fidelity and the target fidelity (that is, the most accurate fidelity, here, TZVP). Generally, it is anticipated that these differences decay monotonically for increasing fidelity. Thirdly and finally, the scatter of the fidelities with respect to the target fidelity is generated to study how these deviate with respect to the target fidelity. The three different preliminary analyses show a systematic ordering of the fidelities giving a nod to the assumed hierarchy as seen in the fidelity difference plots. The fidelity scatter plot also shows a systematic distribution of the energies when compared to the target fidelity of TZVP. In Ref. 13 this was identified as a good preliminary indicator of favorable results in the MFML models. In Figure 2(b), the learning curves of MFML and o-MFML (with OLS optimization) are depicted for the prediction of SCF energies. The multifidelity learning curves can be understood as follows: the addition of a cheaper fidelity systematically decreases the error (here reported as mean absolute error, MAE) which is reported in units of milli-Hartree. With each cheaper fidelity being added, one notices that the corresponding learning curve from Figure 2(b) has a lower offset. The continuing negative slope indicates that further addition of training samples could decrease the MAE of these models. There is no significant difference in MAE between the MFML and o-MFML models and they both perform similary for the prediction of SCF.
(a) Preliminary analysis of multifidelity structure of the first vertical excitation energies for the o-HBDI molecule. The three different preliminary tests for the hierarchy are performed as prescribed in Ref. 13.

### v2

@@ Methods
To build the CheMFi dataset from the WS22 database, 15,000 geometries were sampled from the original 120,000 geometries for each of these molecules. For each of the nine molecules from the WS22 database, 15,000 geometries were evenly sampled from the original 120,000 geometries. To achieve this, every $8^{\rm th}$ geometry for each molecule was selected from the WS22 database resulting in a total of $9\times 15,000=135,000$ point geometries for CheMFi. An even sampling of the original dataset ensures that there are sufficient geometries from all conformations of the molecule. Once these geometries were sampled, they were used to perform point calculations for the QC properties.
All QC calculations were performed with the ORCA(5.0.1) QC package [26]. From these calculations, a diverse set of QC properties were extracted including information of the vertical excitation states such as energies and oscillator strengths. QC calculations were performed at the TD-DFT level of theory with the CAM-B3LYP functional. For each geometry, five fidelities were calculated. These fidelities are the basis set choice of increasing size. In increasing hierarchy of the fidelity, these are: STO-3G, 3-21G, 6-31G, def2-SVP, and def2-TZVP. In the rest of the document, for the most part, these are referred to by their short-hand, i.e., STO3G up to TZVP. The TightSCF keyword was employed to ensure energy convergence of the order of $10^{-9}$ a.u. for each calculation. Resolution of Identity approximation (RIJCOSX) was employed in order to speed up the excitation energy calculations. For any calculation, the maximum memory usage was limited to 2.0 GB. In practice, the ORCA calculations did not use this amount of memory. A total of 10 vertical excitation energies were calculated with each fidelity for each geometry. The complete workflow of the dataset generation process is pictorially depicted in Figure 1.
The list of available multifidelity properties is given in Table 1. The Cartesian coordinates and atomic numbers are taken from the WS22 database. The SCF ground state energies are reported in Hartree units. The first 10 vertical excitation energies are provided in $\rm cm^{-1}$ with their corresponding oscillator strengths and transition dipole moments (in a.u.). The molecular dipole moments are also a property included in the CheMFi dataset with both the nuclear and electronic contributions being separately cataloged in atomic units (a.u.). The rotational spectrum data is also included in the form of Rotational Constants (in $\rm cm^{-1}$) and the total molecular dipole moments (in a.u.) aligned along rotational axes.
As an important contribution to the multifidelity research, the average time to run the QC computations are also provided for each molecule for each fidelity. This information can be further used to benchmark multifidelity models across QC properties as was shown in ref. [15] for excitation energies of arenes. The notion here is to assess the error of a model with respect to the time it takes to generate a training set for that specific model. Such an analysis of the model, as shown in ref. [15], shows the true time benefit of multifidelity models, in this specific case, for MFML. In the CheMFi dataset, the time to run an ORCA calculation for a specific fidelity is provided for each molecule in units of seconds. The parallelization of numerical codes is a challenge unto itself and usually comes with artifacts that arise due to the form of parallelization. This results in non-uniform calculations schemes across basis set sizes and molecules. In order to enforce consistency in the calculation times, ORCA calculations were run on a single compute core for 10-evenly sampled geometries of each molecule. The calculation times as returned by the ORCA software are then averaged over these 10 geometries and reported for each fidelity for each molecule. Thus, the time for a single-core calculation of each fidelity for each molecule is provided in units of seconds to benchmark the time-benefit of multifidelity models against single fidelity models.
Machine Learning Methods
Since this data publication is made along with scripts that enable benchmarking of multifidelity methods for ML-QC, a short digression is made here to discuss the ML methods used in these scripts. The choice of ML method is kernel ridge regression (KRR). The use of diverse ML methods for QC is out of scope of this data publication. The accompanying benchmarking scripts use the MFML and o-MFML methods as described in ref. [14, 15, 16]. These two methods are also used to provide technical benchmarks for the CheMFi dataset including time-benefit of the multifidelity methods over single fidelity methods. Thus, MFML and o-MFML also briefly explained here for the sake of completeness.
Kernel Ridge Regression
Let $\mathcal{T}:=\{(\boldsymbol{X}_{i},E_{i})\}_{i=1}^{{N}_{\rm train}}$ be a training set of size ${N}_{\rm train}$ for molecular descriptors $\boldsymbol{X}_{i}$ and their corresponding QC property, $E_{i}$, be given. A KRR model for the prediction of the QC property $E_{i}$, for an unseen query descriptor $\boldsymbol{X}_{q}$, is denoted by
$P_{\rm KRR}\left(\boldsymbol{X}_{q}\right):=\sum_{i=1}^{N_{\rm train}}\alpha_{i}k\left(\boldsymbol{X}_{q},\boldsymbol{X}_{i}\right)\leavevmode\nobreak\ ,$ (1)
where $k$ is the kernel function. The various coefficients, $\boldsymbol{\alpha}$, are trained by solving $(\boldsymbol{K}+\lambda\boldsymbol{I})\boldsymbol{\alpha}=\boldsymbol{E}$, where $\boldsymbol{K}=\left(k(\boldsymbol{X}_{i},\boldsymbol{X}_{j})\right)_{i,j=1}^{{N}_{\rm train}}$ is the kernel matrix, $\boldsymbol{I}$ the identity matrix, $\boldsymbol{E}=\left(E_{1},E_{2},\ldots,E_{{N}_{\rm train}}\right)^{T}$ is the vector of QC properties from the training set and $\lambda$ a regularization parameter. This work uses two kernels while executing KRR. The first is the Matérn Kernel of first order with the discrete L-2 norm:
$k\left(\boldsymbol{X}_{i},\boldsymbol{X}_{j}\right)=\exp{\left(-\frac{\sqrt{3}}{\sigma}\left\lVert\boldsymbol{X}_{i}-\boldsymbol{X}_{j}\right\rVert_{2}^{2}\right)}\cdot\left(1+\frac{\sqrt{3}}{\sigma}\left\lVert\boldsymbol{X}_{i}-\boldsymbol{X}_{j}\right\rVert_{2}^{2}\right)\leavevmode\nobreak\ ,$ (2)
where $\sigma$ denotes a length scale hyperparameter that determines the width of the kernel. The second kernel used is the Laplacian kernel which is given as
$k\left(\boldsymbol{X}_{i},\boldsymbol{X}_{j}\right)=\exp{\left(-\frac{\left\lVert\boldsymbol{X}_{i}-\boldsymbol{X}_{j}\right\rVert_{1}}{\sigma}\right)}\leavevmode\nobreak\ .$ (3)
Multifidelity Machine Learning
Let the training set for data at some fidelity $f$ be $\mathcal{T}^{(f)}:=\left\{\left(\boldsymbol{X}^{(f)}_{i},y_{i}^{(f)}\right)\right\}_{i=1}^{N^{(f)}}$. Let an ordered hierarchy of fidelities be given as $f\in\{1,2,\ldots F-1,F\}$ where $F$ is the most accurate fidelity and is called the target fidelity. In ref. [14, 16] it was shown that a MFML model can be built:
$P_{\rm MFML}^{(F,\eta_{F};f_{b})}\left(\boldsymbol{X}_{q}\right):=\sum_{\boldsymbol{s}\in\mathcal{S}^{(F,\eta_{F};f_{b})}}\beta_{\boldsymbol{s}}P^{(\boldsymbol{s})}_{\rm KRR}\left(\boldsymbol{X}_{q}\right)\leavevmode\nobreak\ ,$ (4)
where, $f_{b}<F$ is called the baseline fidelity, that is the cheapest fidelity used in the multifidelity scheme. The linear combination is performed for the various sub-models built with different fidelities. Each sub-model is denoted by $\boldsymbol{s}=(f,\eta_{f})$, where $2^{\eta_{f}}=N_{\rm train}^{(f)}$. That is, a sub-model is identified by the fidelity and the number of training samples used at that fidelity. The sub-models are chosen based on a scheme discussed extensively in ref. [16] and form the set of sub-models given by $\mathcal{S}^{(F,\eta_{F};f_{b})}$. The different $\beta_{\boldsymbol{s}}$ are the coefficients of the linear combination. Based on previous research [14], the $\beta_{\boldsymbol{s}}$, are set in for MFML as follows:
$\beta_{\boldsymbol{s}}^{\rm MFML}=\begin{cases}+1,&\text{if }f+\eta_{f}=F+\eta_{F}\\ -1,&\text{otherwise}\end{cases}\leavevmode\nobreak\ .$ (5)
Optimized MFML
The optimized MFML (o-MFML) method was introduced in ref. [16] where the combination of the different sub-models built on different fidelities is optimized on a holdout validation set of reference properties. Such an o-MFML was shown to improve the accuracy of prediction for atomization energies and excitation energies across diverse molecules. o-MFML rewrites Eq. (4) by considering values of the coefficients different from that given in Eq. (5). This is done by optimizing the coefficients on a holdout validation set. In this sense, the different $\beta_{\boldsymbol{s}}$ are treated as hyperparameters and this entire process is akin to a hyper-parameter optimization routine which is commonly used in ML methods.
One can define the validation set for such optimization as $\mathcal{V}^{F}_{\rm val}:=\{(\boldsymbol{X}_{q}^{\rm val},E^{\rm val}_{q})\}_{q=1}^{N_{\rm val}}$. This validation-test set approach is common in ML techniques. The hyperparameters are optimized on the validation set and the model is evaluated on the test set.
For a target fidelity $F$, with $N^{(F)}=2^{\eta_{F}}$ training samples at the target fidelity, for a baseline fidelity $f_{b}$, the o-MFML model is defined as
$P_{\rm o-MFML}^{\left(F,\eta_{F};f_{b}\right)}\left(\boldsymbol{X}_{q}\right):=\sum_{\boldsymbol{s}\in mathcal{S}^{(F,\eta_{F};f_{b})}}\beta_{\boldsymbol{s}}^{\rm opt}\cdot P^{(\boldsymbol{s})}_{\rm KRR}\left(\boldsymbol{X}_{q}\right)\leavevmode\nobreak\ ,$ (6)
where $\beta_{\boldsymbol{s}}^{\rm opt}$ are optimized coefficients. This form of optimization is carried out by solving the following optimization problem:
$\beta_{\boldsymbol{s}}^{\rm opt}=\arg\min_{\beta_{\boldsymbol{s}}}\left\lVert\sum_{v=1}^{N_{\rm val}}\left(y_{v}^{\rm ref}-\sum_{\boldsymbol{s}\in S^{\prime}}\beta_{\boldsymbol{s}}\cdot P^{(\boldsymbol{s})}_{\rm KRR}\left(\boldsymbol{X}_{v}\right)\right)\right\rVert_{p}$
where one minimizes some $p$-norm on $\mathcal{V}^{F}_{\rm val}$. Based on ref. [16], the ordinary least squares (OLS) method is used to optimize the coefficients which implicitly uses a 2-norm.
Data Records
The various QC properties of the CheMFi dataset are stored in separate NumPy (v 1.26.4) npz files for each molecule. These npz files have a dictionary-like format allowing for each property to be accessed via its corresponding key denoted in Table 1. Each property itself is stored as a NumPy ndarray with the first dimension being 15,000 corresponding to the number of geometries. Thus, the QC properties can be accessed by querying the right ID. For example, the SCF ground state energies can be accessed with the key ’SCF’ returning a NumPy ndarray of size $15,000\times 5$ where the second dimension of the array corresponds to the five fidelities used. Similarly, one can access the QC computation times using the key ’t’ which results in a NumPy array of shape (5,)(5,)( 5 , ) corresponding to the five fidelities used. The compute times are stored in units of seconds. An example script to accessing the QC properties is shown in Listing 1.
The dataset itself is hosted on Zenodo at https://zenodo.org/records/11636903 with a detailed README file documenting the key aspects of the data [22]. The README also provides information on how to access the different properties using Python. For the CheMFi dataset, the various scripts involved in generating the data, including ORCA input files and shell scripts to extract properties from the ORCA log files, are stored in the code repository that can be accessed at https://github.com/SM4DA/CheMFi. In addition to these scripts, the code repository also contains Python scripts to perform multifidelity benchmarks on this dataset. These can be launched using the CLI and are a handy tool in setting benchmarks for this dataset using current state of art multifidelity methods.
Technical Validation
In order to verify that this form of sampling did in fact evenly cover the conformation space of each molecule, Uniform Manifold Approximation and Projections (UMAPs) [27] are studied herein. To further validate the CheMFi dataset and its use in benchmarking multifidelity methods, The MFML and o-MFML models prescribed in ref. [16] were tested in predicting ground state energies and the first vertical excitation energies. The multifidelity models are built for different baseline fidelities, which refers to the cheapest fidelity included in the model. For example, a baseline fidelity $f_{b}=$631G implies that the multifidelity model is built up of the fidelities 631G, SVP, and TZVP [16]. In addition to benchmarking the multifidelity models on properties of individual molecules, the models are also tested on using data from all the molecules of the dataset. For this purpose, the ground state energy of all molecules are used to train one single MFML and o-MFML model. This is then tested on predicting the ground state energies of all molecules.
While these broad tests serve as a benchmark for multifidelity models on this dataset, the benchmarks of the other properties and molecules are not reported here. However, it is to be pointed out that the scripts provided can be readily used to generate benchmarks for these cases using standard ML methods such as learning curves. All learning curves are reported for a 10-run average, that is, for 10 random shuffling of the training set as directed in ref. [16].
Conformation Space Coverage
In order to ensure that the 15,000 geometries that are sampled from the WS22 database do cover all of the conformation space as spanned by the complete 120,000 geometries, a short study is performed. UMAPs are a powerful tool to perform dimensionality reduction of data. In addition to this, they are useful tools to visualize the multidimensional feature space in ML [27]. To assess the chemical space coverage achieved with this form of sampling, UMAPs of the molecules can be studied. UMAPs are a dimensionality reduction method that is useful to visualize multidimensional data such as the molecular representations used in ML for QC. The resulting 2D embedding can then be used to visualize the complexity of the conformation space based on the coverage that is observed. To this end, molecular descriptors were first generated for the various molecules. The choice of descriptors for this case was the unsorted Coulomb Matrix (CM) which is calculated as:
$C_{i,j}:=\begin{cases}\frac{Z_{i}^{2.4}}{2}\leavevmode\nobreak\ ,&i=j\\ \frac{Z_{i}\cdot Z_{j}}{\left\lVert\boldsymbol{R}_{i}-\boldsymbol{R}_{j}\right\rVert}\leavevmode\nobreak\ ,&i\neq j\leavevmode\nobreak\ ,\end{cases}$ (7)
where, $Z_{i}$ is the atomic charge of the $i^{\rm th}$ atom of the molecule and $\boldsymbol{R}_{i}$ is its Cartesian coordinate. For this proxy of the conformation space that geometries cover, a 2D UMAP was generated and the resulting plots are shown in Figure 2 for all 9 molecules. It is observed that the UMAPs for CheMFi uniformly cover the conformation space spanned by UMAPs of WS22. From this plot it becomes clear that the geometries sampled for the CheMFi dataset from the WS22 database uniformly cover the entire chemical space of WS22. This is true even in cases of multiple localized clusters as seen in the case of alanine or urocanic acid. Therefore, it becomes evident that even though only 15,000 geometries are sampled from the WS22 database, these do uniformly cover the conformation space of WS22 and should therefore offer the same level of chemical complexity for ML models.
Single Molecule Benchmarks
The technical validation carried out individually for SMA and o-HBDI molecules is in line with the experimental set-up of ref. [15]. A total of 12,288 training samples were chosen to build the multifidelity models. 712 samples were set aside as a validation set for the o-MFML model from ref. [16], and the remaining 2,000 samples were used as a test set. The accuracy of the models are gauged with mean absolute error (MAE) in the form of learning curves. Learning curves display the MAE with respect to increasing number of training samples, here, at the highest fidelity, that is, TZVP. In addition, a special kind of learning curve as seen in ref. [15] are also shown. These are MAEs versus the total time to generate the training set for the MFML model. These special learning curves provide a better picture of the time-benefit of using MFML over conventional single fidelity methods.
Figure 3 reports the multifidelity benchmarking process for SCF ground state energies for SMA. The SLATM molecular descriptor [28] was used with a Laplacian kernel to perform kernel ridge regression. First, the preliminary analyses as recommended by ref. [15] are shown in Figure 3(a). The first of these is to study the distribution of the multifidelity data. The second analysis is the study of mean absolute differences between each fidelity and the target fidelity (that is, the most accurate fidelity, here, TZVP). Generally, it is anticipated that these differences decay monotonically for increasing fidelity. Thirdly and finally, a scatter plot of the energies of the different fidelities with respect to the energies of the target fidelity is generated to study how these deviate with respect to the target fidelity. The three different preliminary analyses show a systematic ordering of the fidelities which confirm the assumed hierarchy as seen in the fidelity difference plots. The fidelity scatter plot also shows a systematic distribution of the energies when compared to the target fidelity of TZVP. In ref. [15] this was identified as a good preliminary indicator of favorable results in the MFML models. In Figure 3(b), the learning curves of MFML and o-MFML (with OLS optimization) are depicted for the prediction of the ground state energies. The multifidelity learning curves can be understood as follows: the addition of a cheaper fidelity systematically decreases the error (here reported as mean absolute error, MAE) which is reported in units of milli-Hartree. With each cheaper fidelity being added, one notices that the corresponding learning curve from Figure 3(b) has a lower offset. The continuing negative slope indicates that further addition of training samples could decrease the MAE of these models. There is no significant difference in MAE between the MFML and o-MFML models and they both perform similarly for the prediction of ground state energies. As a final assessment for SMA, an MAE versus time to generate the MFML training data is also shown in Figure 3(c) for the MFML and o-MFML models. The total time cost for a given MFML (or o-MFML) model is the total time taken to generate all the training samples at the different fidelities. In other words, if one picks the model $P_{\rm MFML}^{(TZVP;631G)}$ with $N_{\rm train}^{TZVP}=2$, then the total time to generate the training set for this model would be given as $T_{\rm MFML}=2\times t_{TZVP}+4\times t_{SVP}+8\times t_{631G}$, where $t_{f}$ is the computation time corresponding to the fidelity $f$. Since the CheMFi dataset comes with the average compute times for each fidelity of each molecule, this allows for a meaningful benchmarking of multifidelity models with this form of analysis. In Figure 3(c) one observes that with addition of cheaper baseline fidelities, the curves show an increased offset along the time axis. Consider the very last data point of the curve corresponding to the reference KRR for o-MFML (right-hand side plot), which corresponds to $\sim 4\cdot 10^{3}$. If one were to draw an imaginary line parallel to the horizontal axis, it would intersect the curve corresponding to the STO3G baseline MFML model around $\sim 10^{3}$. This implies that one could use the o-MFML model with STO3G baseline to achieve similar accuracy as the reference KRR model with a reduction in time cost by a factor of $44\cdot 10^{3}/1\cdot 10^{3}=$ 4. This indeed shows the effectiveness of such forms of multifidelity models over conventional single fidelity methods.
A similar benchmarking procedure was carried out for the prediction of first vertical excitation energies of o-HBDI. In this case, the unsorted CM were used with the Matérn Kernel. As for the case of SMA, a preliminary analysis study was performed with the resulting plots shown in Figure 4(a). The difference in fidelities plot in the center indicates that the assumed hierarchy holds true for the fidelities. However, the fidelity scatter plot on the right hand side shows two distinct clusters. These correspond to the two main conformers of o-HBDI, namely the cis and trans conformers. The scatter plot also shows some cases where the STO3G fidelity covers a wider range of values and is less localized than the other fidelities. This could indicate that the use of STO3G in the MFML models would result in a lower improvement of the accuracy of the model.
The learning curves for the prediction of the first vertical excitation energies of o-HBDI from CheMFi are shown in Figure 4(b). The MAE are reported in $\rm cm^{-1}$ with the axes identically scaled for both MFML and o-MFML. With the addition of cheaper fidelities, the learning curves show a constant reduction in the offset of the MAE as seen in the near parallel learning curves of the different baselines fidelities. As anticipated from the preliminary analysis, the addition of the STO3G fidelity does not provide significant improvement especially for larger training samples. However, this is rectified, as expected, by the o-MFML method which was indeed shown to fix this very issue in ref. [16]. Indeed, in the MAE versus time to generate training data plots seen in Figure 4(c) for o-HBDI, one observes that the STO3G baseline model fails to provide a reasonable improvement in the time benefit unlike observed for the case of SMA in Figure 3(c). However, for models built with with the other baseline fidelities, a time improvement is still visible. These results are a strong indicator towards the possibility of further research in the field of multifidelity methods for QC.
Cumulative use of the dataset
The CheMFi dataset contains multifidelity QC properties of 9 molecules for 15,000 geometries. This totals to $9\times 15,000=135,000$ point calculations of the QC properties. This is therefore the largest collection of multifidelity dataset which can be used in various benchmarking processes. To demonstrate this form of cumulative use of the dataset, multifidelity models from ref. [16] were tested against this in predicting the ground state energies of the molecules. From each molecule of the CheMFi dataset, 1,500 geometries were randomly chosen and compiled into a total of $9\times 1,500=13,500$ data points. From the 13,500 samples, a random set of 11,000 samples were used as the multifidelity training data. Of the remaining, 500 samples were used as a validation set and 2,000 as the holdout test set. With this setup learning curves were generated for the different multifidelity models in the same fashion as prescribed in ref. [16].
The results of this test are shown in Figure 5 for MFML and o-MFML models. The learning curves show a decreasing slope for both cases for the different baseline fidelities. The addition of each cheaper fidelity results in a lower offset of MAE. The constant slope on the log-log axis indicates that addition of training samples can further decrease the MAE. On the right-hand side plot the scatter of reference TZVP versus MFML predicted SCF ground state energies are delineated. Across the energy ranges the MFML model predicts the SCF ground state energies accurately as can be inferred from the scatter of the values being close to the identity mapping line.
For completeness, the MAE versus time to generate the multifidelity training data are reported in Figure 6 for MFML and o-MFML for this cumulative use of the CheMFi dataset. In this specific case for the cumulative use of the CheMFi dataset, the time to generate the multifidelity training data is calculated based on the molecular geometry that is included in the model. It is observed in Figure 6, that the addition of each baseline fidelity results in a distinct reduction in the time cost of generating training data. The resulting curves indicate a time benefit factor of about 6 for the STO3G baseline fidelity.
Usage Notes
In addition to the multifidelity dataset, various tools to assess and benchmark multifidelity methods are also provided. These include scripts to perform preliminary analysis of the data based on the property of choice as recommended in ref. [15], and the scripts to produce learning curves. Further, scripts to generate unsorted Coulomb Matrices, and the global SLATM descriptors are provided which are built upon the qmlcode package [29]. The scripts are easy to use and well documented allowing for a streamlined benchmarking process with an example shown in Listing 2 for the preliminary analysis. Listing 3 shows an example to generate the multifidelity learning curves.
Code availability
Acknowledgements
The authors acknowledge support by the DFG through projects ZA 1175/3-1 as well as through the DFG Priority Program SPP 2363 on “Utilization and Development of Machine Learning for Molecular Applications – Molecular Machine Learning” through the project ZA 1175_4-1. The computations of the quantum chemical properties were carried out on the PLEIADES cluster at the University of Wuppertal, which was supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). The authors would also like to thank the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal.
Author contributions statement
V.V.: Conceptualization, data collection, methodology, validation, writing - original draft preparation, review & editing.
P.Z.: Funding acquisition, supervision, validation, writing - review & editing
Competing interests
Figures & Tables
Figure 1: The workflow of generating the CheMFi dataset by sampling from the WS22 database. 15,000 geometries are used for each molecule resulting in a total of 135,000 single point geometries. For each of these, multiple QC properties are calculated at DFT level of theory with varying basis set sizes to create the diverse multifidelity dataset.
@@ Acknowledgements
Ground State Energies (SCF) (15000, 5)/hE ‘SCF’
Vertical Excitation Energies (15000,5,10)/$\rm cm^{-1}$ ‘EV’
@@
QC Calculation Times (5,)/seconds ‘t’
Table 1: List of properties available in the CheMFi dataset. The corresponding dimension(s) and units of the properties are also given with the npz file key. From the WS22 database [9, 10]
@@
9
10#load QC compute times
11QC_time = data[’t’]
@@
Figure 2: Scatter plots of UMAPs for the various molecules which compose the WS22 database. The UMAPs were generated for the unsorted Coulomb Matrix (CM) molecular descriptor for each molecule. The legend key indicates the geometries which are part of the WS22 and the CheMFi dataset respectively. For all molecules, it can be observed that the CheMFi dataset traverses the entirety of the configuration space that WS22 also covers.
(a) Preliminary analysis of multifidelity structure of SCF ground state energies for the SMA molecule. The three different preliminary tests for the hierarchy are performed as prescribed in ref. [15]. The ground state energies show a normal distribution centered around 0 hE. The difference in the fidelity energies is monotonically decreasing for increasing fidelities indicating that the assumed hierarchy holds. The scatter plot of the energies of different fidelities with respect to the TZVP fidelity show a compact distribution for the most part. With STO3G there is a wider deviation from the identity map (dashed black line).
(b) Learning curves for MFML and o-MFML for the SCF ground state energies of SMA as recorded in the CheMFi database. The reference single-fidelity KRR is also shown by training on TZVP only. The Laplacian kernel was used with a kernel width of 200.0 and regularization of $10^{-10}$. The Global SLATM [28] molecular descriptors were used.
(c) Time versus MAE plots for MFML and o-MFML models predicting the SCF ground state energies of the SMA molecule. The time to generate the training set for MFML is a comprehensive measure of the cost of a multifidelity model as prescribed in ref. [15].
@@
(a) Preliminary analysis of multifidelity structure of the first vertical excitation energies for the o-HBDI molecule. The three different preliminary tests for the hierarchy are performed as prescribed in ref. [15]. The different fidelities show distinct peaks around 0 $\rm cm^{-1}$ with a small bump around $18000\rm cm^{-1}$. The difference in the fidelities shows a distinct hierarchy with reducing difference with increasing fidelity. The scatter plot of energies of different fidelities with respect to the TZVP fidelity show them well distributed except for the STO3G fidelity which shows a wider spread.
@@
(c) Time versus MAE plots for the prediction of first vertical excitation energy for the o-hbdi molecule. The time to generate the MFML training set is calculated as described in ref. [15].